OpenStudy (anonymous):

what 2 fractions are equivalent to 35/56

OpenStudy (anonymous):

What goes into 35 & 56?

OpenStudy (anonymous):

35 / 7 = 5 56 / 7 = 8 5/8 is equivalent.

OpenStudy (anonymous):

You can go bigger also...35 * 2 = 70 56 *2 = 112 so 70/112 is equivalent

OpenStudy (cwrw238):

to simplify fractions first find prime factors: 35 = 5 * 7 56 = 2 * 2 * 2 * 7 take out the 7's and you get 5 / 8